OpenStudy (anonymous):

how do I show that y=(ln(x) + C)/x is a solution to (x^2)y'+xy=1 ?

OpenStudy (anonymous):

find y' and plug it in the equation

OpenStudy (anonymous):

well, y'=1/x^2-y/x, so it would be y'=1/x^2-(ln(x)+C)/x^2.. but I'm getting lost there

OpenStudy (experimentx):

C = yx - ln(x) replace C by that value in y'= ....C

OpenStudy (anonymous):

I forgot that I had to plug in something for C as well, thanks!